# Helios: An extremely low power event-based gesture recognition for always-on smart eyewear

### v2

@@ 1 Introduction
This paper describes Helios: a low-power, real-time, event-based hand gesture recognition system for smart eyewear. Events from a 3mm×$\times$4mm/20mW event camera are passed to a convolutional neural network running on NXP Nano UltraLite which achieves significant power and speed advantages. Helios consumes less than 350mW and has a latency of 60ms from the start of the gesture. In contrast to Meta’s [9], we do not use frame cameras, and our gesture recognition does not work on top of hand skeletal tracking. Helios’s novelty is in being able to go from event streams to gesture recognition directly. The closest to our work is [2] which also uses events for gesture recognition. They however use the DVS-128 camera which is 40mm×$\times$60mm and unsuitable for use with smart eyerwear. A key feature of event-based gesture recognition to work for smart eyewear is robustness to false positives with ego-motion. We expect the user to be walking around while using the smart eyewear. Helios models adversarial classes along with relevant microgestures so that our gesture recognition system does not get triggered by random noise due to human movement. This in contrast to [1, 2, 9] where this problem is not actively addressed.
@@ 3 Machine Vision System
The machine vision system was designed and optimised to enable the Human Machine Interface (HMI) discussed in Section 4 to work with a glasses mounted mono event camera setup. The complexity of doing a full integration into a pair of smart frames or glasses led to an architecture consisting of two parts connected over USB2.0. The first was an event camera USB peripheral consisting of the Prophesee GenX320 sensor and the Cypress CX3 MIPI CSI2 - USB bridge, the event stream was then transmitted over USB2.0 to the second part, a NXP iMX 8M Nano UltraLite compute platform. This arrangement is shown in Figure 2. The compute platform was chosen so it aligned with current smart glass optimised mobile processors, for example the AR1 from Qualcomm. The following sections go into details for each of these components.
@@ 3.3 System assembly
The event camera module and main PCB were assembled into a enclosure designed to fit onto the Meta Ray-Ban glasses. The camera is oriented such that the camera’s optical axis is pointing downward with a small $20^{o}$ rotation around the camera’s $X$-axis (user is facing the Z axis), based on the OpenCV coordinate system, to accommodate an interaction volume further from the body. Finally, the device is connected to the compute platform detailed above and is compatible with any applications built using OpenEB and the MetaVision SDK from Prophesee.
@@ 5.2.5 Model Inference
At inference time the model requires an aggregated TS as the input. To avoid sampling issues where the microgesture occurred during a period that intersects two models inputs, we create an $L$ window-sized aggregated TS after every 80ms period. The intuition for this is that the model will be unsure on predictions where non-complete microgestures have occurred in the input frame, which will be assigned low probabilities by the model. On the other hand, the predicted probability will be higher when the microgesture occurred within the aggregation window. We overcome the issue of the model predicting a microgesture with low confidence by introducing a probability threshold at which a microgesture is deemed to have occurred.
@@ 5.3 Dataset collection for training ML gesture recognition system
Due to the lack of event stream datasets for hand pose estimation or microgesture detection and the difficulty of obtaining accurate ground-truth annotations on real data, we build a highly efficient event stream simulator to generate a large-scale synthetic events dataset with annotations. Our seven classes of interest include: double-pinch, left-swipe, right-swipe, pinch, resting hands, idle and random hand motion. We generated examples of simulated event data per microgesture class for training and validation. The exact number of samples for training and validation sets per gesture class is given in Figure 9. For our testing setup, we use live user testing as described in Section 6.4.
@@ 6.1 Training and Validation Results
In this section, we present the results of our experiments, focusing on the training and validation accuracy, the behavior of the loss functions, and the analysis of the validation confusion matrix. The train and validation datasets are a 90% / 10% split from the simulated dataset.
@@ 6.2.2 Power measurement:
The instantaneous power consumption for the model and event signal processing, measured on an NXP Nano UltraLite, was 340 mW fully utilising two of the four A-53 cores and using 30 MB of RAM. The target instantaneous power is 80 mW, this enables an average power of between 10-20 mW depending on usage, where the system is not running when not in use. To get to this the next step is to quantize the model, based on prior work we estimate that this will reduce the instantaneous power to $\sim$110 mW. The remaining reduction in power will come from architecture refinement and inference execution optimisation, for example using DSP/NPU acceleration.
@@ 6.4.1 Precision and Recall
We evaluate the performance of our ML model with live user testing on 20 participants. For each participant, we collected the results for double-pinch, left microgesture and right microgestures. Each gesture is performed in a randomized order by a user 10 times, thereby leading to 30 gestures per user. The time between each gesture is 1.5s and the time given to complete a gesture is 2s. The precision and recall for 20 users and the learning curve for 5 naive users is given in Table 2. Our recall scores range from 54% to 86%, and precision scores range between 74% to 92%. The reason for double pinch scores to be lower than the swipes are that in the test app we consider the pinch to be distinct from a double pinch, and we do not optimize the TS to be able to cleanly distinguish between them. We expect the double pinch performance to improve if we set the pinch classification threshold to be very high.

### v4

@@ 1 Introduction
This paper describes Helios: a low-power, real-time, event-based hand gesture recognition system for smart eyewear. Events from a 3mm×$\times$4mm/20mW event camera are passed to a convolutional neural network running on NXP Nano UltraLite which achieves significant power and speed advantages. Helios consumes less than 350mW and has a latency of 60ms from the start of the gesture. In contrast to Meta’s [9], we do not use frame cameras, and our gesture recognition does not work on top of hand skeletal tracking. Helios’s novelty is in being able to go from event streams to gesture recognition directly. The closest to our work is [2] which also uses events for gesture recognition. They however use the DVS-128 camera which is 40mm×$\times$60mm and unsuitable for use with smart eyerwear. A key feature of event-based gesture recognition to work for smart eyewear is robustness to false positives with ego-motion. We expect the user to be walking around while using the smart eyewear. Helios models adversarial classes along with relevant microgestures so that our gesture recognition system does not get triggered by random noise due to human movement. This in contrast to [1, 2, 9] where this problem is not actively addressed. Our complete setup is shown in Figure 1.
@@ 3 Machine Vision System
The machine vision system was designed and optimised to enable the Human Machine Interface (HMI) discussed in Section 4 to work with a glasses mounted mono event camera setup. The complexity of doing a full integration into a pair of smart frames or glasses led to an architecture consisting of two parts connected over USB2.0. The first was an event camera USB peripheral consisting of the Prophesee GenX320 sensor and the Cypress CX3 MIPI CSI2 - USB bridge, the event stream was then transmitted over USB2.0 to the second part, a NXP iMX 8M Nano UltraLite compute platform. This arrangement is shown in Figure 2. The compute platform is designed for power-efficient devices using ARM Cortex-A53 cores which aligns with current smart glass optimised mobile processors. The following sections go into details for each of these components.
@@ 3.3 System assembly
The event camera module and main PCB were assembled into a enclosure designed to fit onto the Meta Ray-Ban glasses. The camera is oriented such that the camera’s optical axis is pointing downward with a small $20^{o}$ rotation around the camera’s $X$-axis (user is facing the Z axis), based on the OpenCV coordinate system, to accommodate an interaction volume further from the body. Our smartglasses setup is illustrated in Figure 3. Finally, the device is connected to the compute platform detailed above and is compatible with any applications built using OpenEB and the MetaVision SDK from Prophesee.
@@ 5.2.5 Model Inference
At inference time the model requires an aggregated TS as the input. To avoid sampling issues where the microgesture occurred during a period that intersects two models inputs, we create an $L$ window-sized aggregated TS after every 80ms period. The intuition for this is that the model will be unsure on predictions where non-complete microgestures have occurred in the input frame, which will be assigned low probabilities by the model. On the other hand, the predicted probability will be higher when the microgesture occurred within the aggregation window. We overcome the issue of the model predicting a microgesture with low confidence by introducing a threshold on softmax probabilities, at which a microgesture is deemed to have occurred.
@@ 5.3 Dataset collection for training ML gesture recognition system
Due to the lack of event stream datasets for hand pose estimation or microgesture detection and the difficulty of obtaining accurate ground-truth annotations on real data, we build a highly efficient event stream simulator to generate a large-scale synthetic events dataset with annotations. Our seven classes of interest include: double-pinch (dp), left-swipe (sl), right-swipe (sr), single/moving pinch (mp), resting hands (r), no-hands (nh) and random hand motion (hu). We generated examples of simulated event data per microgesture class for training and validation. The exact number of samples for training and validation sets per gesture class is given in Figure 9. For our testing setup, we use live user testing as described in Section 6.4.
@@ 6.1 Training and Validation Results
In this section, we present the results of our experiments, focusing on the training and validation accuracy, the behavior of the loss functions, and the analysis of the validation confusion matrix. The train and validation datasets are a 90% / 10% split at the sequence from the simulated dataset.
@@ 6.2.2 Power measurement:
The instantaneous power consumption for the model and event signal processing, measured on an NXP Nano UltraLite, was 340 mW fully utilising two of the four A-53 cores and using 30 MB of RAM. This was determined by measuring the current and voltage on the input power rail when the system was idle and running the inference. The difference minus the MIPI-USB camera controller when streaming gives the power consumption of the model. The target instantaneous power is 80 mW, this translate to an average power of approximately 16 mW derived from the number of pickups and session duration from smart watch usage, which is between 1-20% of an 8 hour day [25]. To get to this the next step is to quantize the model, based on prior work we estimate that this will reduce the instantaneous power to $\sim$110 mW. The remaining reduction in power will come from architecture refinement and inference execution optimisation, for example using DSP/NPU acceleration.
@@ 6.4.1 Precision and Recall
We evaluate the performance of our ML model with live user testing on 20 participants. For each participant, we collected the results for double-pinch, left microgesture and right microgestures. Each gesture is performed in a randomized order by a user 10 times, thereby leading to 30 gestures per user. The time between each gesture is 1.5s and the time given to complete a gesture is 2s. The precision and recall for 20 users and the learning curve for 5 naive users is given in Table 2. Our recall scores range from 54% to 86%, and precision scores range between 72% to 92%. The reason for double pinch scores to be lower than the swipes are that in the test app we consider the single pinch to be distinct from a double pinch, and we do not optimize the TS to be able to cleanly distinguish between them. We expect the double pinch performance to improve if we set the single pinch classification threshold to be very high.